In the distribution of succinic acid between ether and water at 15 degrees Celsius 20 ml of the ethereal layer contains 0.092g of acid find the weight of the acid present in 50ml of the aqueous layer in equilibrium with it if the distribution company efficient for succinic acid. Between water and ether is 5.2



1
Expert's answer
2021-11-06T01:32:00-0400

20×0.092×15=5.2×50×m

27.6= 260m

M= 27.6/260

= 0.11g